OpenStudy (andriod09):

fraction problem, too lazy to try to understand what to do, also, hard for me to grasp, Please help!

OpenStudy (andriod09):

\[\frac{1}{x-2}+\frac{3}{x+6}=\frac{1}{2}\]

OpenStudy (anonymous):

Lets's write it like this for now: 1/a + 3/b = 1/2 Now just add up the left side the same way you would do it as if they were numbers. get a lcd, etc...